# Quantum Cosmology, Decoherence, and the Emergence of Classical Spacetime

### v1

@@ IV Coarse Graining and Emergent Classical Dynamics
In this section we evaluate the decoherence functional derived above under two physically motivated coarse-graining prescriptions, following the general logic developed in the cosmological decoherence literature [9, 11, 12]. While the precise numerical rate of decoherence depends on the coarse graining, its qualitative behavior—namely the efficient suppression of interference between macroscopically distinct curvature perturbations—is robust.
IV.1 Horizon-Based Coarse Graining
@@
$N(\eta,\eta^{\prime};k_{L})\simeq\frac{g^{2}}{4}\int_{q>k_{c}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\exp[-2\mathrm{i}q(\eta-\eta^{\prime})]}{q^{2}}.$ (17)
@@
The rapid oscillations render the kernel sharply peaked near $\eta=\eta^{\prime}$, allowing the local approximation
$N(\eta,\eta^{\prime})\simeq\mathcal{N}(\eta)\,\delta(\eta-\eta^{\prime}).$ (18)
With this approximation, the decoherence functional for a frozen superhorizon mode becomes, upon using Eq. (5),
$\Gamma_{k_{L}}\simeq\frac{1}{2}|\Delta\zeta_{k_{L}}|^{2}\int\mathrm{d}\eta\,a^{4}(\eta)\mathcal{N}(\eta).$ (19)
@@
Changing variables to the number of e-folds $N=\ln a$ gives the parametric estimate
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}N}\sim\frac{g^{2}}{H^{2}}\,|\Delta\zeta_{k_{L}}|^{2}\,a^{3}(N),$ (20)
@@
up to numerical factors of order unity. Decoherence therefore accumulates monotonically during inflation, and once the condition $\Gamma_{k_{L}}\gg 1$ is satisfied (cf. Eq. (5)), $\Gamma_{k_{L}}\gg 1$ the reduced density matrix becomes effectively diagonal.
IV.2 EFT-Based Coarse Graining
@@
An alternative prescription is motivated by effective field theory. One introduces a fixed physical cutoff $\Lambda_{\mathrm{phys}}$ satisfying $H\ll\Lambda_{\mathrm{phys}}\ll M_{\mathrm{Pl}}$, and defines the environment as modes with physical momenta above this scale,
$q>a(\eta)\Lambda_{\mathrm{phys}}.$ (21)
Modes traced out in this way never re-enter the system, rendering decoherence effectively irreversible.
@@
For all such modes one has $q/a\gg H$, and the WKB approximation remains valid throughout inflation. The noise kernel again becomes local in time,
$N(\eta,\eta^{\prime})\simeq\mathcal{N}_{\Lambda}\delta(\eta-\eta^{\prime}),$ (22)
with
$\mathcal{N}_{\Lambda}\sim g^{2}\int_{q>a\Lambda_{\mathrm{phys}}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{1}{q^{2}}\sim g^{2}\Lambda_{\mathrm{phys}}a(\eta).$ (23)
The resulting decoherence functional scales as a direct consequence of Eq. (5),
@@
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}N}\sim\frac{g^{2}\Lambda_{\mathrm{phys}}}{H}\ |\Delta\zeta_{k_{L}}|^{2}\ a^{4}(N).$ (24)
Decoherence is thus even more efficient than in the horizon-based prescription and remains insensitive to the details of ultraviolet completion.
@@
In both coarse-graining schemes the essential conclusion is the same: interactions with unobserved degrees of freedom rapidly suppress interference between distinct long-wavelength curvature perturbations, providing a robust dynamical origin for classicality.
IV.3 Order-of-Magnitude Estimates
@@
It is useful to translate the above scaling relations into simple numerical estimates. Observations of the cosmic microwave background imply a dimensionless scalar amplitude $A_{s}\simeq 2.1\times 10^{-9}$, corresponding to a typical root-mean-square curvature perturbation $\zeta_{\rm rms}\sim\sqrt{A_{s}}\sim 5\times 10^{-5}$. Macroscopically distinct semiclassical branches may therefore be characterized by differences $|\Delta\zeta_{k_{L}}|\sim n\,\zeta_{\rm rms}$ with $n\sim 10$100100.
Using the EFT-based coarse graining, the parametric estimate for the decoherence rate yields
@@
$\Gamma_{k_{L}}(\Delta N)\sim\frac{g^{2}}{4}\,\frac{\Lambda_{\mathrm{phys}}}{H}\,|\Delta\zeta_{k_{L}}|^{2}\,\bigl(e^{4\Delta N}-1\bigr),$ (25)
where $\Delta N$ is the number of $e$-folds since horizon exit of the mode $k_{L}$. Taking a representative ratio $\Lambda_{\mathrm{phys}}/H\sim 10$, one finds that decoherence, defined by $\Gamma_{k_{L}}\gtrsim 1$, typically occurs within a few to at most $\mathcal{O}(10)$ $e$-folds for couplings as small as $g\sim 10^{-5}$ and much more rapidly for matter-sector or portal-like couplings $g\sim 10^{-3}$$10^{-1}$. These estimates indicate that environment-induced decoherence is efficient well before the end of inflation for a broad and physically reasonable range of parameters.
V Hartle–Hawking and Tunneling Proposals After Decoherence
@@
V.1 Semiclassical Branch Structure
In the semiclassical regime, the universal wavefunction admits a WKB decomposition of the form
@@ V.1 Semiclassical Branch Structure
where the label $\alpha$ distinguishes different semiclassical branches. For homogeneous minisuperspace variables these branches typically correspond to expanding and contracting classical solutions, while for perturbations they encode different inflationary histories. Importantly, the Wheeler–DeWitt equation is time-reversal invariant, and both types of branches generically appear in the wavefunction for either boundary condition [2, 3].
At the level of the full wavefunction, these branches coexist coherently. The WKB approximation alone does not suppress interference between macroscopically distinct geometries, nor does it select a preferred classical history.
V.2 Reduced Density Matrix in the Branch Basis
The physically relevant object is the reduced density matrix obtained after tracing over unobserved degrees of freedom. Projecting onto the semiclassical branch basis, one finds matrix elements of the schematic form
@@ V.2 Reduced Density Matrix in the Branch Basis
where $\Gamma_{\alpha\beta}$ is the decoherence functional evaluated for the difference between the long-wavelength field configurations associated with branches $\alpha$ and $\beta$.
@@
When $\Gamma_{\alpha\beta}\gg 1$, the off-diagonal elements of the reduced density matrix are exponentially suppressed, and the density matrix becomes approximately diagonal in the branch index. Only at this stage does it become meaningful to interpret the wavefunction as describing an ensemble of mutually exclusive classical histories.
V.3 Expanding versus Contracting Histories
A particularly important application concerns the distinction between expanding and contracting semiclassical branches. In an expanding universe, the available phase space of environmental degrees of freedom grows rapidly with time, leading to strong entanglement between the system and the environment. As shown in Sec. IV, this results in efficient and effectively irreversible decoherence.
In contrast, contracting branches generally generate less entanglement with environmental degrees of freedom. As a result, interference between expanding and contracting branches is suppressed primarily in favor of the expanding histories, which acquire robust classical records. This dynamical asymmetry gives operational content to the intuition often associated with the tunneling proposal, but it arises only after decoherence is taken into account.
V.4 Boundary Conditions versus Classicality
The analysis above clarifies the logically distinct roles played by boundary conditions and decoherence. The Hartle–Hawking and tunneling proposals determine the relative amplitudes $A_{\alpha}$ of semiclassical branches in the universal wavefunction. Decoherence, by contrast, determines which of these branches become classical by suppressing quantum interference.
Statements such as “the tunneling proposal predicts an expanding universe” are therefore incomplete unless supplemented by a discussion of environment-induced decoherence. Without decoherence, expanding and contracting branches remain part of a coherent superposition. With decoherence, expanding branches are dynamically selected as classical due to their stronger coupling to environmental degrees of freedom.
V.5 Emergent Arrow of Time
Neither the no-boundary nor the tunneling proposal imposes a fundamental arrow of time at the level of the Wheeler–DeWitt equation. In the present framework, the cosmological arrow of time emerges dynamically from the growth of entanglement between long-wavelength perturbations and their environment. Decoherence increases monotonically with expansion, rendering the arrow of time an emergent, rather than fundamental, property of classical cosmology.
VI Discussion and Outlook

### v2

@@ III.2 Coarse graining and decoherence rates
In this section we evaluate the decoherence functional derived above under two physically motivated coarse-graining prescriptions, following the general logic developed in the cosmological decoherence literature [10, 12, 13, 18, 19]. While the precise numerical rate of decoherence depends on the coarse graining, its qualitative behavior—namely the efficient suppression of interference between macroscopically distinct curvature perturbations—is robust.
III.3 Horizon-based coarse graining
@@
$N(\eta,\eta^{\prime};k_{L})\simeq\frac{\lambda_{\mathrm{eff}}^{2}}{4}\int_{q>k_{c}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\exp[-2\mathrm{i}q(\eta-\eta^{\prime})]}{q^{2}}.$ (33)
@@
The rapid oscillations render the kernel sharply peaked near $\eta=\eta^{\prime}$, allowing the local approximation $N(\eta,\eta^{\prime})\simeq\mathcal{R}(\eta)\,\delta(\eta-\eta^{\prime})$. The decoherence functional for a frozen superhorizon mode then becomes
$\Gamma_{k_{L}}\simeq\frac{1}{2}\,|\Delta\zeta_{k_{L}}|^{2}\int\mathrm{d}\eta\,a^{4}(\eta)\,\mathcal{R}(\eta).$ (34)
@@
Changing variables to the number of e-folds $\mathcal{N}=\ln a$ gives the parametric estimate
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}\mathcal{N}}\sim\frac{\lambda_{\mathrm{eff}}^{2}}{H^{4}}\,|\Delta\zeta_{k_{L}}|^{2}\,a^{3}(\mathcal{N}),$ (35)
@@
up to numerical factors of order unity. Decoherence therefore accumulates monotonically, and once $\Gamma_{k_{L}}\gg 1$ the reduced density matrix becomes effectively diagonal.
III.4 EFT-based coarse graining
@@
An alternative prescription is motivated by effective field theory. One introduces a fixed physical cutoff $\Lambda_{\mathrm{phys}}$ satisfying $H\ll\Lambda_{\mathrm{phys}}\ll M_{\mathrm{Pl}}$, and defines the environment as modes with physical momenta above this scale, $q>a(\eta)\,\Lambda_{\mathrm{phys}}$. Modes traced out in this way never re-enter the system, rendering decoherence effectively irreversible.
@@
For all such modes $q/a\gg H$, and the WKB approximation remains valid throughout inflation. The noise kernel again becomes local in time, $N(\eta,\eta^{\prime})\simeq\mathcal{R}_{\Lambda}\,\delta(\eta-\eta^{\prime})$, with
$\mathcal{R}_{\Lambda}\sim\lambda_{\mathrm{eff}}^{2}\int_{q>a\Lambda_{\mathrm{phys}}}\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{1}{q^{2}}\sim\lambda_{\mathrm{eff}}^{2}\,\Lambda_{\mathrm{phys}}\,a(\eta),$ (36)
giving
@@
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}\mathcal{N}}\sim\frac{\lambda_{\mathrm{eff}}^{2}\,\Lambda_{\mathrm{phys}}}{H^{5}}\,|\Delta\zeta_{k_{L}}|^{2}\,a^{4}(\mathcal{N}).$ (37)
@@
Decoherence is thus even more efficient than in the horizon-based prescription and remains insensitive to the details of ultraviolet completion. In both schemes, interactions with unobserved degrees of freedom rapidly suppress interference between distinct long-wavelength curvature perturbations, providing a robust dynamical origin for classicality.
III.5 Order-of-magnitude estimates
@@
Observations of the cosmic microwave background imply $A_{s}\simeq 2.1\times 10^{-9}$, giving $\zeta_{\mathrm{rms}}\sim\sqrt{A_{s}}\sim 5\times 10^{-5}$. Macroscopically distinct semiclassical branches may be characterized by $|\Delta\zeta_{k_{L}}|\sim n\,\zeta_{\mathrm{rms}}$ with $n\sim 10$$100$.
Using the EFT-based coarse graining,
@@
$\Gamma_{k_{L}}(\Delta\mathcal{N})\sim\frac{\hat{\lambda}^{2}}{4}\,\frac{\Lambda_{\mathrm{phys}}}{H}\,|\Delta\zeta_{k_{L}}|^{2}\,\bigl(\mathrm{e}^{4\Delta\mathcal{N}}-1\bigr),$ (38)
where $\hat{\lambda}\equiv\lambda_{\mathrm{eff}}/H^{2}$ and $\Delta\mathcal{N}$ counts e-folds since horizon exit. Setting $\Gamma_{k_{L}}(\Delta\mathcal{N}_{\mathrm{dec}})\simeq 1$ gives
$\Delta\mathcal{N}_{\mathrm{dec}}=\frac{1}{4}\ln\!\left[1+\frac{4}{\hat{\lambda}^{2}\,(\Lambda_{\mathrm{phys}}/H)\,|\Delta\zeta_{k_{L}}|^{2}}\right].$ (39)
For $\Lambda_{\mathrm{phys}}/H\sim 10$, the explicit values are given in table 1. In the minimal spectator model, $\hat{\lambda}=\tfrac{3}{2}m_{\sigma}^{2}/H^{2}$, so, for example, $m_{\sigma}/H=0.1$ gives $\hat{\lambda}=0.015$ and $\Delta\mathcal{N}_{\mathrm{dec}}\approx 5.67$ ($n=10$) and $4.52$ ($n=100$); $m_{\sigma}/H=0.3$ gives $\hat{\lambda}=0.135$ and $4.57$ ($n=10$) and $3.42$ ($n=100$). Decoherence occurs within a few to at most $\mathcal{O}(10)$ e-folds across the full range, confirming efficient classicalisation well before the end of inflation.
Table 1: Decoherence time $\Delta\mathcal{N}_{\mathrm{dec}}$ (e-folds after horizon exit) from equation (39), for $\Lambda_{\mathrm{phys}}/H=10$ and branch separation $|\Delta\zeta_{k_{L}}|=n\,\zeta_{\mathrm{rms}}$ with $\zeta_{\mathrm{rms}}\simeq 5\times 10^{-5}$.
$\hat{\lambda}$ $\Delta\mathcal{N}_{\mathrm{dec}}$ $(n=10)$ $\Delta\mathcal{N}_{\mathrm{dec}}$ $(n=100)$
$10^{-5}$ 9.33 8.18
$10^{-3}$ 7.03 5.88
$10^{-1}$ 4.72 3.57
IV Hartle–Hawking and tunneling proposals after decoherence
@@
IV.1 Semiclassical branch structure
In the semiclassical regime, the universal wavefunction admits a WKB decomposition
@@ IV.1 Semiclassical branch structure
where $\alpha$ distinguishes different semiclassical branches. For homogeneous minisuperspace variables these branches typically correspond to expanding and contracting classical solutions, while for perturbations they encode different inflationary histories. The Wheeler–DeWitt equation is time-reversal invariant, and both types of branch generically appear for either boundary condition [1, 2]. These branches coexist coherently at the level of the full wavefunction; the WKB approximation alone does not select a preferred classical history.
IV.2 Reduced density matrix in the branch basis
Projecting the reduced density matrix onto the semiclassical branch basis yields matrix elements of the schematic form
@@ IV.2 Reduced density matrix in the branch basis
where $\Gamma_{\alpha\beta}$ is the decoherence functional for the difference between the long-wavelength field configurations associated with branches $\alpha$ and $\beta$. When $\Gamma_{\alpha\beta}\gg 1$, the off-diagonal elements are exponentially suppressed and the density matrix becomes approximately diagonal. Only at this stage does it become meaningful to interpret the wavefunction as an ensemble of mutually exclusive classical histories.
IV.3 Expanding versus contracting histories
The reduced density matrix element between the two WKB branches can be written as
$\rho_{+-}(a)=A_{+}(a)\,A_{-}^{*}(a)\,\mathrm{e}^{\mathrm{i}M(S_{+}(a)-S_{-}(a))}\,\mathcal{D}(a),$ (42)
where
$\mathcal{D}(a)\equiv\mathrm{Tr}_{\mathrm{env}}\!\left[U_{+}(a)\,\rho_{\mathrm{env}}\,U_{-}^{\dagger}(a)\right],$ (43)
where $U_{\pm}$ denote the environmental evolution operators conditioned on the expanding and contracting WKB backgrounds, and $M$ is the large semiclassical parameter multiplying the gravitational Hamilton–Jacobi action. We use $\mathcal{D}$ for the branch-overlap factor throughout this section to distinguish it from the dissipation kernel $D$ in (4). For Gaussian environmental states the decoherence factor factorises over modes,
$\mathcal{D}(a)=\prod_{k\in\mathrm{env}}\mathcal{D}_{k}(a)=\exp\!\left[-\Gamma_{+-}(a)+\mathrm{i}\Theta_{+-}(a)\right],$ (44)
with $\Gamma_{+-}(a)=-\sum_{k\in\mathrm{env}}\ln|\mathcal{D}_{k}(a)|$. Once $\Gamma_{+-}\gg 1$, interference between the two semiclassical geometries is exponentially suppressed.
For an environmental oscillator mode whose conditional states on the two branches are Gaussian with kernels $\Omega_{k}^{(+)}$ and $\Omega_{k}^{(-)}$,
$|\mathcal{D}_{k}|=\left[\frac{4\,\mathrm{Re}\,\Omega_{k}^{(+)}\,\mathrm{Re}\,\Omega_{k}^{(-)}}{|\Omega_{k}^{(+)}+\Omega_{k}^{(-)\,*}|^{2}}\right]^{1/4},$ (45)
so branch decoherence is controlled by how differently the two backgrounds squeeze the same environmental mode. We now evaluate $|\mathcal{D}_{k}|$ explicitly using the Bunch–Davies mode functions already established in section III.
Gaussian kernels on expanding and contracting backgrounds
The vacuum state on each branch is Gaussian: $\Psi^{\pm}_{k}[\phi_{k}]\propto\exp(-\frac{1}{2}\Omega_{k}^{(\pm)}|\phi_{k}|^{2})$, with the kernel related to the mode function by
$\Omega_{k}^{(\pm)}=\mathrm{i}\,\frac{\partial_{\eta}u_{k}^{(\pm)}}{u_{k}^{(\pm)}}.$ (46)
On the expanding branch ($a=-1/(H\eta)$, $\eta<0$), the Bunch–Davies solution is $u_{k}^{(+)}\propto(-\eta)^{3/2}H^{(1)}_{\nu}(-k\eta)$. On the contracting branch ($a=+1/(H\eta)$, $\eta>0$), the corresponding vacuum is $u_{k}^{(-)}\propto\eta^{3/2}H^{(2)}_{\nu}(k\eta)$. Using the Hankel recurrence $(H^{(1)}_{\nu})^{\prime}(z)=H^{(1)}_{\nu-1}(z)-(\nu/z)H^{(1)}_{\nu}(z)$ with $z=-k\eta$, equation (46) gives
$\Omega_{k}^{(+)}$ $=-\mathrm{i}k\!\left[\frac{\Delta}{z}+\frac{H^{(1)}_{\nu-1}(z)}{H^{(1)}_{\nu}(z)}\right],$ (47)
$\Omega_{k}^{(-)}$ $=+\mathrm{i}k\!\left[\frac{\Delta}{z^{\prime}}+\frac{H^{(2)}_{\nu-1}(z^{\prime})}{H^{(2)}_{\nu}(z^{\prime})}\right],$ (48)
where $z^{\prime}=k\eta$ and $\Delta=3/2-\nu$. The sign reversal reflects the opposite orientation of conformal time on the two branches. Throughout, $z=-k\eta>0$ is the dimensionless wavenumber on the expanding branch ($\eta<0$), while $z^{\prime}=k\eta>0$ is its counterpart on the contracting branch ($\eta>0$); they are numerically equal at equal $|k\eta|$, so the comparison in equations (49)–(50) is made at equal $|z|$.
Real and imaginary parts
Writing $H^{(1)}_{\nu}=J_{\nu}+\mathrm{i}Y_{\nu}$, defining $\mathcal{W}(z)\equiv J_{\nu}^{2}(z)+Y_{\nu}^{2}(z)>0$, and using the standard cross-order Wronskian (Abramowitz & Stegun 9.1.16) $J_{\nu}Y_{\nu-1}-J_{\nu-1}Y_{\nu}=2/(\pi z)$, one finds
$\mathrm{Re}\,\Omega_{k}^{(+)}$ $=+\frac{2k}{\pi z\,\mathcal{W}(z)},$ (49)
$\mathrm{Im}\,\Omega_{k}^{(+)}$ $=-k\!\left[\frac{\Delta}{z}+\frac{J_{\nu-1}J_{\nu}+Y_{\nu-1}Y_{\nu}}{\mathcal{W}(z)}\right].$ (50)
Note that $\mathrm{Re}\,\Omega_{k}^{(+)}>0$ for all $z$, as required for the Gaussian wavefunction $\Psi^{+}_{k}\propto\exp(-\tfrac{1}{2}\Omega_{k}^{(+)}|\phi_{k}|^{2})$ to be normalizable. Since $H^{(2)}_{\nu}=(H^{(1)}_{\nu})^{*}$, the contracting branch satisfies $\mathrm{Re}\,\Omega_{k}^{(-)}=\mathrm{Re}\,\Omega_{k}^{(+)}$ and $\mathrm{Im}\,\Omega_{k}^{(-)}=-\mathrm{Im}\,\Omega_{k}^{(+)}$ at equal $|k\eta|$.
Closed-form result
Setting $R\equiv\mathrm{Re}\,\Omega_{k}^{(+)}$ and $I\equiv\mathrm{Im}\,\Omega_{k}^{(+)}$, and noting $\Omega_{k}^{(+)}+\Omega_{k}^{(-)\,*}=2R+2\mathrm{i}I$, substitution into equation (45) gives
$\boxed{|\mathcal{D}_{k}(z)|=\left[\frac{R^{2}}{R^{2}+I^{2}}\right]^{1/4}=\frac{1}{(1+\tan^{2}\!\theta_{k})^{1/4}},}$ (51)
where $\theta_{k}\equiv\arg\Omega_{k}^{(+)}$ is the squeezing angle of the expanding-branch vacuum. The overlap depends only on the expanding-branch state; the contracting branch enters through the symmetry $\mathrm{Im}\,\Omega^{(-)}=-\mathrm{Im}\,\Omega^{(+)}$.
Massless case ($\nu=3/2$, exact).— For $\nu=3/2$ the half-integer Bessel functions give $\mathcal{W}=(2/\pi z)(1+z^{-2})$. Using $H^{(1)}_{1/2}(z)=-i\sqrt{2/(\pi z)}e^{iz}$ and $H^{(1)}_{3/2}(z)=-\sqrt{2/(\pi z)}e^{iz}(1+i/z)$ one finds $H^{(1)}_{1/2}/H^{(1)}_{3/2}=iz/(z+i)$, so equation (47) gives $\Omega_{k}^{(+)}=-ik\cdot iz/(z+i)=kz(z-i)/(z^{2}+1)$, yielding
$R=+\frac{kz^{2}}{z^{2}+1}>0,\quad I=-\frac{kz}{z^{2}+1},$ (52)
and the exact closed form
$|\mathcal{D}_{k}(z)|\big|_{\nu=3/2}=\left[\frac{z^{2}}{z^{2}+1}\right]^{1/4}.$ (53)
The intermediate step is immediate: $R^{2}/(R^{2}+I^{2})=(k^{2}z^{4}/(z^{2}+1)^{2})/[(k^{2}z^{4}+k^{2}z^{2})/(z^{2}+1)^{2}]=z^{2}/(z^{2}+1)$, confirming the closed form. This satisfies $|\mathcal{D}_{k}|\to 1$ as $z\to\infty$ (subhorizon) and $|\mathcal{D}_{k}|\sim z^{1/2}\to 0$ as $z\to 0$ (superhorizon).
Massive fields ($\nu<3/2$).— Superhorizon asymptotics give $\mathcal{W}\sim Y_{\nu}^{2}\propto z^{-2\nu}$, so $R\propto z^{2\nu}\to 0$ while $I$ remains finite, giving
$|\mathcal{D}_{k}(z)|\sim z^{\nu}\to 0\quad(z\to 0),$ (54)
verified numerically to better than $0.3\%$ for $\nu=1.20$, $1.35$, $1.45$ (table 2).
Table 2: Numerical verification of the superhorizon power law $|\mathcal{D}_{k}|\propto z^{\nu}$. The ratio $|\mathcal{D}_{k}(z_{1})|/|\mathcal{D}_{k}(z_{2})|$ at $z_{1}=0.003$, $z_{2}=0.010$ is compared with $(z_{1}/z_{2})^{\nu}$.
$\nu$ $m_{\sigma}/H$ Numerical ratio $(z_{1}/z_{2})^{\nu}$
1.45 0.55 0.17470 0.17451
1.35 0.95 0.19694 0.19684
1.20 1.32 0.23590 0.23580
The behavior of $|\mathcal{D}_{k}(z)|$ is illustrated in figure 2.
Figure 2: Branch-overlap factor $|\mathcal{D}_{k}(z)|$ as a function of $z=-k\eta$, from subhorizon ($z\gg 1$) to superhorizon ($z\to 0$), for four spectator masses. Curves are labelled by spectator mass $m_{\sigma}/H$ (equivalently $\nu$); the massless case ($\nu=3/2$) uses the exact closed form (53); massive curves use the general formula (51) with Bessel functions evaluated numerically. All curves approach unity subhorizon and decay to zero superhorizon. The dotted vertical line marks horizon crossing ($z=1$).
Accumulated branch-decoherence functional
Converting the mode sum to a phase-space integral,
$\Gamma_{+-}(\mathcal{N})=\frac{1}{2\pi^{2}}\int_{0}^{k_{\max}}\mathrm{d}k\,k^{2}\left(-\ln|\mathcal{D}_{k}(z_{k}(\mathcal{N}))|\right),$ (55)
with $z_{k}(\mathcal{N})=k/(H\,\mathrm{e}^{\mathcal{N}})$. Numerical evaluation using equations (49)–(53) shows that $\Gamma_{+-}$ crosses unity within $\mathcal{N}\approx 0.5$ e-folds for all values of $m_{\sigma}/H$ considered, and reaches $\mathcal{O}(10^{3})$ within four to five e-folds (figure 3). Branch separation is therefore not only efficient but effectively irreversible.
Two complementary decoherence rates.— It is important to distinguish the two accumulated functionals that appear in this paper, which measure related but conceptually distinct quantities. Figure 3 shows $\Gamma_{+-}(\mathcal{N})$ computed directly from the geometric branch-overlap formula (55), $\Gamma_{+-}=-\sum_{k}\ln|\mathcal{D}_{k}|$, using the Bunch–Davies mode functions. This quantity is independent of the spectator coupling $\hat{\lambda}$: it measures how differently the two semiclassical backgrounds squeeze each environmental mode, and it crosses unity within $\approx 0.5$ e-folds for all spectator masses considered. The coupling-dependent noise-kernel estimate of section III.2, by contrast, depends explicitly on $\hat{\lambda}$ and on the branch-separation amplitude $|\Delta\zeta_{k_{L}}|$; for $\hat{\lambda}=10^{-3}$ the scalar threshold is crossed at $\mathcal{N}\approx 7$ e-folds (table 1). The two calculations are complementary: equation (55) isolates the geometric asymmetry between branches encoded in the Bunch–Davies mode functions, providing a lower bound on the e-fold at which geometry alone distinguishes the branches, while the noise-kernel estimate quantifies the rate at which system–environment interactions of a specific spectator sector exploit that asymmetry to suppress off-diagonal density-matrix elements. Both calculations confirm irreversible branch separation well within inflation. In brief: the geometric $\Gamma_{+-}$ measures the orthogonality of two distinct semiclassical geometries — a purely kinematic consequence of the Hilbert-space structure — established almost instantly upon horizon crossing, while the noise-kernel estimate measures the rate at which a specific perturbation history is recorded by a spectator environment to a precision of $\zeta_{\mathrm{rms}}$, which requires the sustained accumulation of environmental records over several e-folds.
Figure 3: Accumulated branch-decoherence functional $\Gamma_{+-}(\mathcal{N})$ on linear (left) and logarithmic (right) scales. The dotted horizontal line marks $\Gamma_{+-}=1$. The threshold is crossed within $\approx 0.5$ e-folds for all spectator masses; thereafter $\Gamma_{+-}$ grows rapidly, confirming irreversible branch separation. The massless curve uses the exact closed form (53); massive curves use the general formula (51). This geometric $\Gamma_{+-}$ is independent of the coupling $\hat{\lambda}$ and provides a lower bound on the decoherence timescale; the coupling-dependent noise-kernel estimate yields a later crossing at $\mathcal{N}\approx 7$ e-folds for $\hat{\lambda}=10^{-3}$ (table 1).
Physical interpretation
Equation (51) makes the mechanism transparent. The overlap is governed by the squeezing angle $\theta_{k}=\arg\Omega_{k}^{(+)}$ of the expanding-branch vacuum. With $R>0$ and $I<0$ (equation (49)), one has $\tan\theta_{k}=I/R=-1/z$, so $\theta_{k}\in(-\pi/2,0)$ for all $z>0$. Subhorizon modes have $z\gg 1$, giving $|\theta_{k}|\approx 0$ and $|\mathcal{D}_{k}|\approx 1$ — the two branches are indistinguishable. As a mode crosses the horizon, inflationary particle production drives $z\to 0$, so $|\theta_{k}|\to\pi/2$ and $R=kz^{2}/(z^{2}+1)\to 0$, giving $|\mathcal{D}_{k}|\to 0$. On the contracting branch the mode function $u_{k}^{(-)}$ does not undergo the same superhorizon amplification, so the squeezing angle on that branch remains $\mathcal{O}(1)$; the two-branch overlap therefore vanishes because of the asymmetry between $\Omega_{k}^{(+)}$ and $\Omega_{k}^{(-)}$. The asymmetry is a direct consequence of the asymmetry in superhorizon mode production — the same physical process that generates the observed scale-invariant curvature spectrum. Branch decoherence and the generation of classical primordial perturbations are two facets of the same inflationary squeezing. The same asymmetry between expanding and contracting branches applies in bouncing cosmology scenarios where a Wheeler–DeWitt contracting phase precedes inflation [25].
IV.4 Relation to previous work
Polarski and Starobinsky emphasized the semiclassical appearance of cosmological perturbations and the role of decoherence in rendering squeezed states effectively classical [10]. Kiefer and collaborators developed the reduced-density-matrix viewpoint in quantum cosmology [7, 13, 17]. Burgess, Holman, and Hoover analyzed the decoherence of inflationary fluctuations from the perspective of effective field theory and subhorizon environmental interactions [12], with subsequent work extending the analysis to the open-EFT framework [19]. The quantum discord and einselection approach to inflationary classicality has been developed in [23, 18].
The main point of the present analysis is not to propose a wholly new decoherence mechanism, but to assemble a more explicit bridge between ingredients that are often discussed separately: the origin of the effective long-mode coupling to spectator sectors, the coarse-graining dependence of the resulting decoherence exponent, and the branch-overlap criterion relevant for Hartle–Hawking and tunneling wavefunctions.
IV.5 Boundary conditions versus classicality
The Hartle–Hawking and tunneling proposals determine the relative amplitudes $A_{\alpha}$ of semiclassical branches. Decoherence determines when the overlap between those branches becomes negligibly small. Without decoherence, expanding and contracting branches remain part of a coherent superposition; with decoherence, the branch-overlap factor $\mathcal{D}(a)$ in (42) determines when the two histories can be treated as an effectively classical ensemble. The explicit calculation in section IV establishes that for the inflationary coarse graining adopted here, the expanding branch is the one for which $|\mathcal{D}_{k}|\to 0$ as modes freeze out superhorizon, while the contracting branch retains $|\mathcal{D}_{k}|\sim\mathcal{O}(1)$. This is a derived result, not an inference: it follows directly from the asymmetry in superhorizon mode production encoded in equations (51) and (53).
IV.6 Emergent arrow of time
Neither boundary proposal imposes a fundamental arrow of time at the level of the Wheeler–DeWitt equation. The origin of the cosmological arrow of time has long been connected to the special low-entropy initial conditions of the universe [26]; the present framework reframes this: the explicit evaluation of $\Gamma_{+-}(\mathcal{N})$ in figure 3 shows that the cosmological arrow of time emerges from the same inflationary squeezing that produces the primordial perturbation spectrum: each mode that crosses the horizon drives the squeezing angle $\theta_{k}\to\pi/2$ and contributes irreversibly to $\Gamma_{+-}$. The expanding branch becomes classically distinguishable from the contracting branch within the first e-fold of inflation, and the asymmetry grows without bound thereafter. Decoherence therefore supplies the mechanism by which one orientation of semiclassical evolution becomes stably classical, without inserting a fundamental arrow by hand.
V Discussion and outlook